Inventory flow assumptions Arrow, Inc. uses a perpetual inventory system. On January 22, 2018, the company had 200 units of a particular product on hand, with a total cost of $2,400. The per-unit costs were Date Purchase Unit Cost Total Cost tit Ending inventory, 2017 Jan. 10 purchase Total on hand 50 150 200 S 9 S13 S 450 1950 $2,400 On January 24, 2018, Arrow sold 65 units of this product. Using the three flow assumptions listed below, compute (1) the cost of goods sold, and (2) the cost of the inventory of this product on hand after this sale. Show your computations as per below format (a) Average cost Cost of goods sold Inventory remaining after sale LIFO Cost of goods sold Inventory remaining after sale (c FIFO Cost of goods sold Inventory remaining after sale
Answer #1

(a) Average Cost

(1) Cost of Goods Sold - 65 x 2,400/200 = 780

(2) Inventory Remaining after Sale = (2,400- 780) = 1,620

(b) LIFO

(1) Cost of Goods Sold - 65 x 13 = 845

(2) Inventory Remaining after Sale = (2,400- 845) = 1,555

(C) FIFO

(1) Cost of Goods Sold - (50 x 9) + ( 15 x 13) = 645

(2) Inventory Remaining after Sale = (2,400- 645) = 1,755